OpenStudy (anonymous):

A company that manufactures cardboard boxes has employed you to find the dimensions of the box that minimizes the cost to make each box. The boxes are to have a square base and a volume of 1200 cubic inches. In addition, the top and bottom of the box cost twice as much as the sides.

OpenStudy (ahsome):



OpenStudy (ahsome):

Since it has a square bottom, the sides must be the same

OpenStudy (ahsome):

\[h\times s^2=1200\]

OpenStudy (ahsome):

\[s^2 = 2h\times s\]In terms of cost

OpenStudy (ahsome):

\[s2=2(h×s)\]